is it possible that P(A)=17/3 ?
Found 3 solutions by Alan3354, ewatrrr, mathmate:
Answer by Alan3354(69443)

is it possible that P(A)=17/3 ?
-----------
No.
P(A) <= 1.

Answer by ewatrrr(24785)


Good Morning!
is it possible that P(A)=17/3 = 5 2/3 ? Nope!
Note: 0 ≤ P(A) ≤ 1

Answer by mathmate(429)


Question:
is it possible that P(A)=17/3 ?

Solution:
If P(A)=17/3 is meant to say that the probability of event A happening is 17/3, then the answer is it is not possible.
By Kolmogorov's axioms of probability, probability of ANY event must lie between zero and one, inclusively.